OpenStudy (aravindg):

In a triangle sin A ,sin B ,sin C are in A P then... a) altitudes are in AP B)the al;titudes are in HP C)the altitudes are in G P D)none of these

OpenStudy (aravindg):

@phi, @myininaya

OpenStudy (aravindg):

@amistre64

OpenStudy (apoorvk):

I would guess B. Harmonic progression. since due to Sine law, inverse relations build up between the sides and sine values. I haven't solved it properly, but I ll let you know.

OpenStudy (aravindg):

@Mani_Jha

OpenStudy (aravindg):

@satellite73

OpenStudy (aravindg):

@KingGeorge

OpenStudy (anonymous):

ok at first glance i am going to say arithmetic. let me see if can draw a picture

OpenStudy (anonymous):

nope that is not working out for me

OpenStudy (aravindg):

hmm ..these days my qns makes people here to think more

OpenStudy (aravindg):

:P

OpenStudy (anonymous):

yes it is a lot of thinking. so far i have thought this \[h_1=b\sin(A)\] \[h_2=c\sin(B) = c(\sin(A)+d)=c\sin(A)+cd\] \[h_3=a\sin(C)=a(\sin(A)+2d)=a\sin(A)+2ad\]

OpenStudy (anonymous):

and \[\frac{c}{\sin(C)}=\frac{a}{\sin(A)}\] \[\frac{c}{\sin(A)+2d}=\frac{a}{\sin(A)}\] so \[c=\frac{a(\sin(A)+2d}{\sin(A)}\] and now maybe we can write \(h_2\) in term so only of "a" and \(\sin(A)\) and "d"

OpenStudy (aravindg):

@nikvist

OpenStudy (anonymous):

\[h_2=\frac{a(\sin(A)+2d)}{\sin(A)}(\sin(A)+d)\]}

OpenStudy (anonymous):

\[h_3=a(\sin(A)+2d)\] so it looks like \(h_2\) is a multiplye of \(h_3\)

OpenStudy (anonymous):

similarly \[b=\frac{a(\sin(A)+d)}{\sin(A)}\] so \[h_1=\frac{a(\sin(A)+d)}{\sin(A)}\times \sin(A)=a(\sin(A)+d)\]

OpenStudy (anonymous):

did i screw this up yes?

OpenStudy (anonymous):

*yet

OpenStudy (aravindg):

i think there is an easier way

OpenStudy (anonymous):

certainly a more systematic way i am sure

OpenStudy (anonymous):

you have \[\sin(A), \sin(B)=\sin(A)+d, \sin(C)=\sin(A)+2d\] then they are in arithemtic progression now you have to write the altitudes in term of the sides and the sines

OpenStudy (aravindg):

@FoolForMath , @Hero

OpenStudy (anonymous):

but the only way i can see to do it is write \[h_1=b\sin(A)\] and then solve for "b" in terms of "a" and \(\sin(A)\)

OpenStudy (aravindg):

@Sarkar

OpenStudy (experimentx):

I am trying to find those three angles programmatically ... lol

OpenStudy (anonymous):

@AravindG I think you should give them a medal, don't be stingy.... at least they're trying to help you!

OpenStudy (experimentx):

improved my code by at lot still no luck ... anyone here good at programming??

OpenStudy (aravindg):

@kreshnik u see i was having dinner .. no qn i have left without giving medals!!